OpenStudy (leahhhmorgannn):

How can you use the process of factoring and completing the square in a quadratic function to show zeros, extreme values, and symmetry of the graph and interpret these in terms of a context? I've got the zeros part. The rest just... I'm not sure.

OpenStudy (leahhhmorgannn):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

how would you complete the square on something like x^2 + 5x + 6 ?

OpenStudy (leahhhmorgannn):

I'm not really sure.

jimthompson5910 (jim_thompson5910):

x^2 + 5x + 6 is the same as 1x^2 + 5x + 6

jimthompson5910 (jim_thompson5910):

1x^2 + 5x + 6 is in the form ax^2 + bx + c a = 1 b = 5 c = 6

jimthompson5910 (jim_thompson5910):

plug a = 1 and b = 5 into h = -b/(2a) and tell me what you get for h

OpenStudy (leahhhmorgannn):

well h = -5/(2 times 1) h = -5 (2) h = -10? I'm not really great at stuff like that, which is why I'm basically not understanding any algebra.

jimthompson5910 (jim_thompson5910):

it would be -5/2 = -2.5

OpenStudy (leahhhmorgannn):

mm, okay. I thought I'd learned that you'd multiply something like that, but okay.

jimthompson5910 (jim_thompson5910):

so h = -2.5

jimthompson5910 (jim_thompson5910):

we plug this into x^2 + 5x + 6 to find the value of k

jimthompson5910 (jim_thompson5910):

k = h^2 + 5h + 6 k = (-2.5)^2 + 5(-2.5) + 6 k = -0.25

jimthompson5910 (jim_thompson5910):

so we know this a = 1 (given) h = -2.5 k = -0.25 and we plug all that into y = a(x-h)^2 + k to get this y = a(x-h)^2 + k y = 1(x-(-2.5))^2 + (-0.25) y = (x+2.5)^2 - 0.25

jimthompson5910 (jim_thompson5910):

so y = x^2 + 5x + 6 turns into y = (x+2.5)^2 - 0.25 after we complete the square

jimthompson5910 (jim_thompson5910):

what does completing the square get us? it tells us what the vertex and axis of symmetry are vertex: (h,k) = (-2.5, -0.25) axis of symmetry: x = h = -2.5

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):